The chords of contact of the pair of tangents drawn from each point on the line 2x + y = 4 to the circle x2+y2 = 1 pass through a fixed point-
 
        (a)    (2,4)
 
(b)    (-1/2,-1/4)
 
(c)    (1/2, 1/4)
 
(d)    (-2, -4)

Grade:12th pass

2 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

Equation of chord of contact:
hx+ky=1
(h, k) lie on the line, we have
2h+k=4
hx+(4-2h)y=4
=>(1/2, 1/4)

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Equation of chord of contact:
hx+ky=1
(h, k) lie on the line, we have
2h+k=4
hx+(4-2h)y=4
=>(1/2, 1/4)

Thanks and Regards